OpenStudy (anonymous):

The volume of a box is represented by the function V(x) = x^3 + 6x^2 + 11x + 6. The height of the box is x + 2. If the area of the base is 24cm^2, determine the height of the box.

OpenStudy (zpupster):

the equation for V=lwh so we know lw=24 h=x+2 V= x^3 + 6x^2 + 11x + 6 our equation now looks like this; x^3 + 6x^2 + 11x + 6=(x+2)(24) divide both sides by x+2

OpenStudy (zpupster):

what do you get??

OpenStudy (anonymous):

I got x^2 + 4x + 3 on the left side and R = 0/x+2

OpenStudy (anonymous):

On the right side, (x + 2)(24)... (x+2)(24) = 24x+48. I have to divide 24x+48 by (x+2)??? I am back to original if I divide it though..

OpenStudy (zpupster):

ok x^2 + 4x + 3=24 -->(x+2) cancel on rt

OpenStudy (anonymous):

Yes

OpenStudy (zpupster):

now subtract 24 from both sides and factor the quadratic equation

OpenStudy (anonymous):

ohhhhhhh! Alright

OpenStudy (zpupster):

what did you get

OpenStudy (anonymous):

I am still solving!

OpenStudy (anonymous):

(x-3) and (x+7) ???

OpenStudy (zpupster):

set both those to zero and solve ignore the neg. answer

OpenStudy (anonymous):

Okay, give me a sec

OpenStudy (zpupster):

x-3=0 x+7=0

OpenStudy (anonymous):

Yes, I am doing that currently!

OpenStudy (anonymous):

Does this seem right? (x-3) is height.

OpenStudy (zpupster):

x-3=0 x=3 so 3+2=5 is height

OpenStudy (zpupster):

plug 3 in and check your answer

OpenStudy (anonymous):

Alright, Thank you!

OpenStudy (zpupster):

good job!!

OpenStudy (anonymous):

:D Thank you so much!

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.